# What Changed and What Could Have Changed? State-Change Counterfactuals for Procedure-Aware Video Representation Learning

### v4

@@ 4.1 Implementation Details
Text Generation. To generate all state-change descriptions and state-change counterfactuals, we use Llama 3.2 8B [17] for efficiency. We incorporate off-the-shelf text-encoder FLAVA [70] to pre-extract features of the annotated narrations and summaries and our generated text descriptions.
@@
Temporal Action Segmentation. We evaluate representations on temporal action segmentation, where models output the frame-wise action label for untrimmed, long videos, requiring the ability to reason about procedures and recognize the start and end of state transformations. We evaluate various representations by adopting the popular temporal action segmentation model ASFormer [95] that takes video features as input. We follow prior work [39] for the same training strategy and evaluation protocol and metrics including F1 scores, Edit distance, and frame-wise Accuracy, on the cooking procedure datasets, Georgia Tech Egocentric Activities (GTEA) [20] and EgoPER [37] Dataset. GTEA consists of 28 egocentric instructional videos capturing daily kitchen activities at 15 frames per second. EgoPER contains 368 videos spanning 28 hours of cooking scenarios.
@@
Error Detection. We evaluate representations on unsupervised error detection on the EgoPER dataset [37] where models are trained with error-free videos and output labels for each frame on whether it is an error during testing time. The error categories include step omission, step addition, step modification, step slip, and step correction. We follow [37] and use its EgoPED model [37], and equip it with different video representations. Specifically, the training process of EgoPED consists of two stages: action segmentation and prototype learning. In stage 1, the temporal action segmentation model, ASFormer, is trained to predict frame-wise action labels on the error-free set by taking pre-extracted video features as input. In stage 2, the model learns prototype features for each class of action step, e.g., “Take bowl from microwave,”, with contrastive step prototype learning [54, 37]. For each class of activity, multiple prototypes are extracted for error detection at inference. Note that each activity is trained independently, such as Make Coffee. During the inference, EgoPED calculates the similarity score of observed frame features and any learned prototypes. A threshold is further used to determine whether the observed test frame is erroneous.
@@
Action Retrieval & Recognition. To demonstrate the applicability of the learned representations in the core video understanding tasks, we evaluate zero-shot multi-instance retrieval on Epic-Kitchens (EK) [12] and zero-shot action recognition on Charades-Ego (CE) [69]. Following prior work [42, 2], both tasks are implemented via text-to-vision alignment by selecting the textual label with the highest similarity score to the given visual sample.
@@
Table 2: Error detection results on the EgoPER dataset. The underline denotes the second best. HTM denotes the HowTo100M dataset [49]. Text denotes the VLM model with a trainable text encoder.
@@
Action Phase Classification & Frame Retrieval. To assess more fine-grained and short-term procedure awareness, we test representations on action phase classification and zero-shot frame retrieval on the Align-Ego-Exo (AE2) dataset [91]. The AE2 dataset contains egocentric and exocentric videos displaying the actions Break Eggs, Pour Milk, Pour Liquid, and Tennis Forehand. Each action is divided into two to four phases [91]. We follow [91] and train an SVM to predict the per-frame action phase labels for the classification task and use nearest neighbors for the zero-shot retrieval task. Note that we merge the validation and test sets for more robust results.
@@ 4.3 Error Detection Results
Moreover, we achieve competitive performance on the AUC metric. Note that although we follow prior work [37] to include frame-wise metric AUC, this metric may not be effective in measuring the performance and the procedure awareness of representations. As discussed in EgoPER [37], the heuristic random method can achieve competitive results in AUC but fail to localize erroneous segments as shown in the major metric EDA. This is also true when comparing I3D and earlier methods, such as HF2-VAD [46] and S3R [86]. We hope our extensive experimental results can inspire future work to study more effective evaluation metrics for error detection. In summary, our representation learning method achieves state-of-the-art performance in EDA and competitive performance in AUC, highlighting the strong effectiveness of counterfactual reasoning in procedure awareness.
Table 3: Multi-Instance Retrieval (MIR) on Epic-Kitchens (EK) and Zero-shot Action Recognition on Charades-Ego (CE). We report PT ckpt mAP for the Action Recognition task (see [2]).
@@ 4.4 Action Retrieval & Recognition Results
Table 3 presents benchmark results on the action retrieval and recognition. We found existing procedure-aware representations perform significantly worse on short-term action recognition tasks. Yet, our representations trained with both short- and long-term state-changes and counterfactuals demonstrate strong applicability to the popular video recognition tasks and datasets.
@@ 4.5 Action Phase Classification & Retrieval Results
Table 4 shows that our model outperforms others on average across actions on both the ego+exo and ego settings, demonstrating its effectiveness in short-term procedures and strong view-point generalization. Interestingly, CLIP [56] demonstrates strong performance. We conjecture this is due to the focus on short-term action in its image-to-text alignment. Full results with individual actions are available in Section B.1 in the supplementary material.
@@
[…]
@@ 4.6 Ablation Study
In Table 5, we study the effectiveness of state changes (Before and After, state-change counterfactuals (SC-CF), and video-level state-change counterfactuals (Missing-step CF and Misordered-step CF) by ablating each of them and evaluating on temporal action segmentation with GTEA and error detection with EgoPER. Note that ID 1 can be seen as Weak HierVL as it is trained with only the video encoder. In addition, since Weak HierVL only uses narrations from other videos in a batch as negatives, it can be seen as a random description approach. All ablations (including Ours) are pre-trained with a batch size of 12 on each GPU.
Temporal Action Segmentation. One interesting insight on GTEA is that using only clip-level state changes and their counterfactuals (ID 3) achieves competitive performance, outperforming most other ablated models. Also, using only one type of video-level counterfactual (ID 4 and 5) may mislead the model to overfit a certain type of counterfactual, resulting in lower performance. By considering all the proposed states and counterfactuals together, we show that they complement each other, thus achieving the best performance.
Error Detection. The results on EgoPER show that our proposed state changes and their counterfactuals can improve performance by combining them all together, verifying the effectiveness of recognizing erroneous labels. In addition, we observe interesting results showing that with only clip-level counterfactuals (ID 3), the model performs slightly worse than its baseline (ID 2). We conjecture this is because the model overfits short-term procedures and overlook long-term activities procedures and their counterfactual reasoning.
@@ 5 Conclusions
In this work, we present a novel procedure-aware video representation learning framework that first incorporates state-change descriptions and state-change counterfactuals in clip-level alignment, enhancing causal reasoning of action transformations. Then, it utilizes video-level counterfactuals that perturb the local actions and create hypothesized scenarios to facilitate the understanding of activity procedures. Our learned representations demonstrate strong effectiveness in terms of procedure awareness and achieve state-of-the-art results on several benchmarks. For future work, a promising direction for further strengthening procedure awareness for video representation learning with state-change counterfactuals is to generate multi-modal counterfactual examples for procedure-aware activities, such as synthetic long-form videos showcasing mistakes and the corresponding corrections.
@@ 6 Acknowledgment
This work was supported in part by the National Science Foundation under award DRL-2112635 to the AI Institute for Engaged Learning, and he National Science and Technology Council (NSTC) under grants 113-2628-E-A49 -022 - and 114-2628-E-A49 -007 -, the Higher Education Sprout Project of the National Yang Ming Chiao Tung University, and Ministry of Education (MOE), the Yushan Fellow Program Administrative Support Grant.
@@ Appendix A Text Description Generation
In this section, we present the text-generation process used in this work. We generate clip-level state-change descriptions, i.e., Before, After and State-change counterfactual; and video-level state-change counterfactuals, i.e., Missing-key-Step Counterfactuals and Misordered-Step Counterfactuals. In this work, we use Llama 3.1 [17], the latest version of Llama at the time of implementation. To generate the text descriptions for the significantly large dataset Ego4D [23], we select Llamma3.1 8B for efficiency.
@@ A.1 Prompt Design
We feed the clip narrations and video summaries to Llamma to generate the corresponding states and counterfactuals. Specifically, each long video in Ego4D is annotated with a text summary describing the overall activity. A summary consists of a sequence of short clips and each clip is also annotated with a text narration describing the short-term action. Below, we present the generation of clip- and video-level texts separately:
@@
Given a clip’s narration, $t_{i}$, we prompt by first feeding the context implemented by Llama:
@@
Missing-key-Step Counterfactuals and Misordered-Step Counterfactuals
For video-level text generation, we feed the whole sequence of clip narrations $t_{0},...,t_{K}$ in a long video, and the video’s summary $S_{i}$. For Missing-key-step Counterfactual, we use the following context prompt:
@@ A.2 Generated Examples - Clip-level
Before: The paper lies flaton a table, surrounded byother papers and office supplies.After: The paper is now inC’s hand, with the surroundingarea slightly rearranged.SC-CF 1: The paper remains flaton the table, untouched.SC-CF 2: A corner of the paperis folded up, but the rest remains flat.SC-CF 3: The paper has been torn inhalf, with one half on the tableand the other half on the floor.
@@ A.4 Quality of Generated Descriptions
We asked 22 human raters and Gemini 2.5 Pro to evaluate 300 and 1000 pairs of LLM-generated state changes (SC) and their counterfactuals (CF), respectively, by Likert-scoring from 1 to 5 for Relevance ($R$) and Plausibility ($P$). Human scores: SCR: 4.95, CFR: 4.84; SCP: 4.73, CFP: 3.87. Gemini scores: SCR: 4.85, CFR: 4.58 ; SCP: 4.91, CFP: 4.57. Despite being reasonable and relevant, we found that the generated CFs occasionally reflect low-probability scenarios, suggesting a tradeoff between creativity and realism in LLMs. Yet, the ablations in Table 5 verify their effectiveness and robustness on procedure-aware tasks.

### v6

@@ 4.1 Implementation Details
Text Generation. To generate all state-change descriptions and state-change counterfactuals, we use Llama 3.1 8B [17] for efficiency. We incorporate off-the-shelf text-encoder FLAVA [69] to pre-extract features of the annotated narrations and summaries and our generated text descriptions.
@@
Temporal Action Segmentation. We evaluate representations on temporal action segmentation, in which the goal is to output frame-wise action labels for untrimmed videos. This requires modeling procedures and recognizing the start and end of state transformations. We evaluate various representations by adopting the popular temporal action segmentation model ASFormer [94] that takes video features as input. We follow prior work [39] for the same training strategy and evaluation protocol and metrics including F1 scores, Edit distance, and frame-wise Accuracy, on the cooking procedure datasets, Georgia Tech Egocentric Activities (GTEA) [20] and EgoPER [37] Dataset. GTEA consists of 28 egocentric instructional videos capturing daily kitchen activities at 15 frames per second. EgoPER contains 368 videos spanning 28 hours of cooking scenarios.
@@
Error Detection. We evaluate representations on unsupervised error detection on the EgoPER dataset [37] where models are trained with error-free videos and output labels for each frame on whether it is an error. The error categories include step omission, step addition, step modification, step slip, and step correction. We follow [37], and use the EgoPED model [37] and equip it with different video representations. Specifically, the training process of EgoPED consists of two stages: action segmentation and prototype learning. In stage 1, the temporal action segmentation model, ASFormer, is trained to predict frame-wise action labels on the error-free set by taking pre-extracted video features as input. In stage 2, the model learns prototype features for each class of action step, e.g., “Take bowl from microwave,” with contrastive step prototype learning [54, 37]. For each class of activity, multiple prototypes are extracted for error detection at inference. Note that each activity is trained independently, such as Make Coffee. During the inference, EgoPED calculates the similarity score of observed frame features and any learned prototypes. A threshold is further used to determine whether the observed test frame is erroneous.
@@
Action Retrieval & Recognition. To demonstrate the applicability of the learned representations in core video understanding tasks, we evaluate zero-shot multi-instance retrieval on Epic-Kitchens (EK) [12] and zero-shot action recognition on Charades-Ego (CE) [68]. Following prior work [42, 2], both tasks are implemented via text-to-vision alignment by selecting the textual label with the highest similarity score to the given visual sample. Note that [42] reports overfitting when applying the model trained on Ego4D to Charades-Ego, and thus uses a task-specific checkpoint. We follow [2] and report the results of our pre-training checkpoint (denoted as PT ckpt in their work) rather than using task-specific checkpoints.
@@
Table 2: Error detection results on the EgoPER dataset. Bold indicates best and underline is second best. HTM denotes the HowTo100M dataset [49]. Text denotes the VLM model with a trainable text encoder.
@@
Action Phase Classification & Frame Retrieval. To assess fine-grained and short-term procedure awareness, we test representations on action phase classification and zero-shot frame retrieval on the Align-Ego-Exo (AE2) dataset [90]. The AE2 dataset contains egocentric and exocentric videos of actions Break Eggs, Pour Milk, Pour Liquid, and Tennis Forehand. Each action is divided into two to four phases [90]. We follow [90] and train an SVM to predict the per-frame action phase labels for the classification task and use nearest neighbors for the zero-shot retrieval task. Note that we merge the validation and test sets for more robust results.
@@ 4.3 Error Detection Results
Moreover, we achieve competitive performance on the AUC metric. Note that although we follow prior work [37] to include frame-wise metric AUC, this metric may not be effective in measuring the performance and the procedure awareness of representations; as discussed in EgoPER [37], a heuristic random method can achieve competitive results in AUC but fails to localize erroneous segments according to EDA. This is also true when comparing I3D and earlier methods, such as HF2-VAD [46] and S3R [85]. We hope our extensive experimental results can inspire future work to study more effective evaluation metrics for error detection. In summary, our representation learning method achieves state-of-the-art performance in EDA and competitive performance in AUC, highlighting the strong effectiveness of counterfactual reasoning in procedure awareness.
Table 3: Multi-Instance Retrieval (MIR) on the Epic-Kitchens (EK) dataset [12] and Zero-shot Action Recognition on the Charades-Ego (CE) dataset [68].
@@ 4.4 Action Retrieval & Recognition Results
Table 3 presents benchmark results on action retrieval and recognition. We found existing procedure-aware representations perform significantly worse on short-term action recognition tasks, while our representations trained with both short- and long-term state-changes and counterfactuals work well on popular video recognition tasks and datasets.
@@ 4.5 Action Phase Classification & Retrieval Results
Table 4 shows that our model outperforms others on average across actions on both the ego+exo and ego settings, demonstrating its effectiveness in short-term procedures and strong view-point generalization. Interestingly, CLIP [56] demonstrates strong performance. We conjecture this is due to the focus on short-term action in its image-to-text alignment. Full results with individual actions are available in the supplementary material.
@@
Table 5: Ablation study involving: Before & After states, State-change Counterfactuals (SC-CF), Missing-step (K-CF) and Misordered (M-CF) on GTEA for temporal action segmentation and EgoPED for error detection. Note that we calculate EDA by averaging all categories of activity. Metrics: Acc. and EDA; mean $\scriptstyle\pm$ sd of 5 and 10 runs, respectively.
@@ 4.6 Ablation Study
In Table 5, we study the effectiveness of state changes (Before and After), state-change counterfactuals (SC-CF), and video-level state-change counterfactuals (Missing-step CF and Misordered-step CF) by ablating each of them and evaluating on temporal action segmentation with GTEA and error detection with EgoPER. Note that the first row (ID 1) can be seen as Weak HierVL as it is trained with only the video encoder. In addition, since Weak HierVL only uses narrations from other videos in a batch as negatives, it can be seen as a random description approach. All ablations (including Ours) are pre-trained with a batch size of 12 on each GPU.
Temporal Action Segmentation. One interesting insight on GTEA is that using only clip-level state changes and their counterfactuals (ID 3) achieves competitive performance, outperforming most other ablated models. Also, using only one type of video-level counterfactual (ID 4 and 5) may mislead the model to overfit to a certain type of counterfactual, resulting in lower performance. By considering all the proposed states and counterfactuals together, we show that they complement each other, thus achieving the best performance.
Error Detection. The results on EgoPER show that our proposed state changes and their counterfactuals can improve performance by combining them all together, verifying the effectiveness of recognizing erroneous labels. In addition, we observe interesting results showing that with only clip-level counterfactuals (ID 3), the model performs slightly worse than its baseline (ID 2). We conjecture this is because the model overfits to short-term procedures and overlooks long-term activity procedures and their counterfactual reasoning.
@@ 5 Conclusions
In this work, we present a novel procedure-aware video representation learning framework that first incorporates state-change descriptions and state-change counterfactuals in clip-level alignment, enhancing causal reasoning of action transformations. Then, it uses video-level counterfactuals that perturb the local actions and create hypothesized scenarios to facilitate the understanding of activity procedures. Our learned representations demonstrate strong effectiveness in terms of procedure awareness and achieve state-of-the-art results on several benchmarks. For future work, a promising direction may be to generate multi-modal counterfactual examples for procedure-aware activities, such as synthetic long-form videos showcasing mistakes and the corresponding corrections.
@@ 6 Acknowledgment
Authors CK, FR, JH, and DC were supported in part by the National Science Foundation under award DRL-2112635 to the AI Institute for Engaged Learning and by the IU Luddy Artificial Intelligence Center. YC was supported in part by the National Science and Technology Council under grants 113-2628-E-A49 -022 - and 114-2628-E-A49 -007 -, by the Higher Education Sprout Project of National Yang Ming Chiao Tung University, and by the Ministry of Education through the Yushan Fellow Program Administrative Support Grant. Computation resources were supported in part by the Lilly Endowment, Inc. through its support to the IU Pervasive Technology Institute.
@@ Appendix A Text Description Generation
In this section, we present the text-generation process used in this work. We generate clip-level state-change descriptions, i.e., Before, After and State-change counterfactual; and video-level state-change counterfactuals, i.e., Missing-Step Counterfactuals and Misordered Counterfactuals. In this work, we use Llama 3.1 [17], the latest version of Llama at the time of implementation. To generate the text descriptions for the significantly large dataset Ego4D [23], we select Llama3.1 8B for efficiency.
@@ A.1 Prompt Design
We feed the clip narrations and video summaries to Llama to generate the corresponding states and counterfactuals. Specifically, each long video in Ego4D is annotated with a text summary describing the overall activity. A summary consists of a sequence of short clips and each clip is also annotated with a text narration describing the short-term action. Below, we present the generation of clip- and video-level texts separately:
@@
Given a clip’s narration, $t_{i}$, we prompt by first feeding the context input into Llama:
@@
Missing-Step Counterfactuals and Misordered Counterfactuals
For video-level text generation, we feed the whole sequence of clip narrations $t_{0},...,t_{K}$ in a long video, and the video’s summary $S_{i}$. For Missing-step Counterfactual, we use the following context prompt:
@@ A.2 Generated Examples - Clip-level
Before: The paper lies flaton a table, surrounded byother papers and office supplies.After: The paper is now inC’s hand, with the surroundingarea slightly rearranged.CF 1: The paper remains flaton the table, untouched.CF 2: A corner of the paperis folded up, but the rest remains flat.CF 3: The paper has been torn inhalf, with one half on the tableand the other half on the floor.
@@ A.4 Quality of Generated Descriptions
We asked 22 human raters and Gemini 2.5 Pro to evaluate 300 and 1000 pairs of LLM-generated state changes (SC) and their counterfactuals (CF), respectively, by Likert-scoring from 1 to 5 for Relevance ($R$) and Plausibility ($P$). Human scores: SCR: 4.95, CFR: 4.84; SCP: 4.73, CFP: 3.87. Gemini scores: SCR: 4.85, CFR: 4.58 ; SCP: 4.91, CFP: 4.57. Despite being reasonable and relevant, we found that the generated CFs occasionally reflect low-probability scenarios, suggesting a tradeoff between creativity and realism in LLMs. Yet, the ablations in the main paper verify their effectiveness and robustness on procedure-aware tasks.